Can you please tell me how to solve this? [100*(x-500)-1350]*14=10,000
Thanks,
Tanya

Found 2 solutions by algebrapro18, stanbon:
Answer by algebrapro18(249)

[100*(x-500)-1,350]*14=10,000
14(100x-50,000-1,350)=10,000
(1400x-718900=10,000)/100
14x-7189 = 100
14x = 7289
x = 520.64 apporximatly

Answer by stanbon(75887)

[100*(x-500)-1350]*14=10,000
[100x-50000-1350]*14=10000
[100x-51350]=714.2857
100x=52064.2857
x=520.6429
Cheers,
Stan H.